If a gallon of gas weighs 6 1/5 pounds and the contents of your gas tank weighs 80 3/5 pounds, how many gallons of gas do you have in your tank?
Answer by stanbon(75887)

If a gallon of gas weighs 6 1/5 pounds and the contents of your gas tank weighs 80 3/5 pounds, how many gallons of gas do you have in your tank?
-----------------------------------
x = (80 3/5)/(6 1/5) = (403/5)/(31/5) = 403/31 = 13 gallons
=============================
Cheers,
Stan H.
======================